Rabulani Ltd recently had better than expected earnings which it does not expect to achieve again.
The company wants to distribute 90% of the earnings available to common shareholders for the year through a share repurchase at the current share price, instead of a paying out a dividend. The buy-back will be offset against retained earnings, which comprises nearly the entirety of the company’s equity. The company currently has 10 000 000 shares outstanding trading at R5 each, R60 000 000 in total assets (including the earnings available to common shareholders for the year), R50 000 000 in total liabilities and earnings available to common shareholders is R10 000 000.
Required:
Determine how many shares will be bought back and how the share buy-back would influence the number of shares outstanding in the market. Also discuss the effect the buy-back will have on remaining shareholders if the future earnings available to common shareholders are expected to be a constant R10 000 000 per year.
1
Expert's answer
2020-10-01T06:53:25-0400

At present 10,000,000 shares outstanding

Earning available =10,000,000

Current market price = 5/-

Amount available for Buy Back =90% of 10,000,000

=9,000,000/-

No. of shares that can be brought back = 9,000,000/5

=1,800,0000 shares

No. of shares remaining after Buy Back

=10,000,000-1,800,000

=8,200,000 shares.

Eject of the Buy Back on EPS

price Buy Back

EPS = 10,000,000/8,200,000

=1.2195 per share.